# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 5 Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
6 Simulation study
The finite-sample properties of the circmax estimators and of the bootstrap approaches have been investigated in a large scale Monte-Carlo simulation study. Two asymptotic regimes were considered: first, the case where the block size $r$ is fixed and the number of disjoint blocks is increasing, and second, the case where $n$ is fixed and the block size is treated as a tuning parameter. All performance metrics in the subsequent paragraphs have been empirically approximated based on $N=5,000$ simulation repetitions.
6.1 Fixed block size
Throughout, we fix the block size $r$, and consider the estimation of target parameters associated with the law of $\bm{M}_{r}$, as well as bootstrap approximations for the respective estimation error. For illustrative purposes, we restrict attention to the one dimensional case and the most vanilla target parameter: the expected value $\mu_{r}=\operatorname{E}[M_{r}]$ of the block maximum distribution. The respective disjoint, sliding and circular block maxima estimators are $\hat{\mu}_{n}^{({\operatorname{mb}})}:=n^{-1}\sum_{i=1}^{n}M_{r,i}^{({\operatorname{mb}})}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, where ${\operatorname{cb}}(k)$ denotes the circmax estimator with parameter $k\in\{2,3\}$. The finite-sample performance is assessed for the following time series model.
Model 6.1 (ARMAX-GPD-Model).
$(X_{t})_{t\in\mathbb{Z}}$ is a stationary real-valued time series whose stationary distribution is the generalized Pareto distribution $\mathrm{GPD}(0,1,\gamma)$ with c.d.f.
$F_{\gamma}(x):=\begin{cases}\left(1-(1+\gamma x)^{-1/\gamma}\right)\mathds{1}(x\geq 0),\quad&\gamma>0,\\ \left(1-(1+\gamma x)^{-1/\gamma}\right)\mathds{1}(0\leq x\leq-1/\gamma),\quad&\gamma<0,\\ \left(1-\exp(-x)\right)\mathds{1}(x\geq 0),&\gamma=0.\end{cases}$
[…]
Note in passing that $(X_{t})_{t}$ from Model 6.1 is exponentially beta mixing and satisfies Condition 2.1 with $Z\sim\mathrm{GEV}(\gamma)$, $a_{r}=\{r(1-\beta)\}^{\gamma}$ and $b_{r}=\{r(1-\beta)^{\gamma}-1\}/\gamma$. Moreover, the mean and variance of $M_{r}$ exist for $\gamma<1$ and $\gamma<1/2$, respectively, and in the latter case an application of Theorems 2.3 and 3.2 shows that, for $r=r_{n}\to\infty$ such that $r=o(n)$ and $\log(n)=o(r^{1/2})$,
$\frac{\sqrt{n/r}}{(r(1-\beta))^{\gamma}}\big{(}\hat{\mu}_{n}^{({\operatorname{mb}})}-\mu_{r}\big{)}\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{{\operatorname{mb}}}),\quad{\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\},$
where the asymptotic variance is given by
$\sigma^{2}_{{\operatorname{db}}}=\begin{cases}\frac{g_{2}-g_{1}^{2}}{\gamma^{2}},&\gamma<1/2,\gamma\neq 0,\\ \frac{\pi^{2}}{6},&\gamma=0,\end{cases}\qquad\sigma^{2}_{{\operatorname{cb}}(k)}=\sigma^{2}_{{\operatorname{sb}}}=\begin{cases}4\Gamma(-2\gamma)I(\gamma),&\gamma<0,\\ 4(\log 4-1),&\gamma=0,\\ -\frac{g_{2}}{\gamma}I(\gamma)&0<\gamma<1/2.\end{cases}$ (6.2)
Here, $\Gamma$ denote the gamma function, $g_{j}=\Gamma(1-j\gamma)$ and $I(\gamma):=2\int_{0}^{1/2}\{\alpha_{2\gamma}(w)-1\}\{w^{-\gamma-1}(1-w)^{-\gamma-1}\}\,\mathrm{d}w$ with $\alpha_{c}(w)=w^{-1}\int_{0}^{w}(1-z)^{c}\,\mathrm{d}z$. Details are provided in Section A.
For the simulation experiments, the block size was fixed to $r=90$ as it roughly corresponds to the number of days within a season - a common block size in environmental applications. The effective sample size (i.e., the number of disjoint blocks of size $r$, abbreviated by $m=\lfloor n/r\rfloor$ hereafter) has been varied between $50$ and $100$ resulting in total sample sizes $n$ between $4,500$ and $9,000$
Performance of the estimators. We start by comparing the four estimators in terms of their variance, squared bias and mean squared error (MSE). For assessing the bias, the true first moment $\mu_{r}$ was approximated by a preliminary Monte Carlo simulation based on $10^{6}$ independent block maxima from Model 6.1 with block size $r=90.$ The results are summarized in Figure 2. As suggested by the theory, both the sliding blocks estimator and the two circular blocks estimators perform uniformly better than the respective disjoint blocks estimator. The improvement gets smaller as the observations are getting more heavy-tailed; a phenomenon that has already been observed in the literature; see e.g. Bücher and Zanger, (2023), Bücher and Staud, (2024). It is noteworthy that the additional bias effect introduced in the circular blocks estimators does not contribute in a relevant way; its contribution to the overall MSE was found to be at most $0.2\%$ and often much smaller. Furthermore, the serial dependence does not change the qualitative results significantly.
Figure 2: Mean estimation with fixed block size $r=90$. Top three rows: Mean squared error $\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{mb}})})$. Bottom three rows: relative MSE with respect to the disjoint blocks method, i.e., $\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{mb}})})/\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{db}})})$.
Performance of the bootstrap. We consider each of the bootstrap estimators $\hat{\mu}_{n}^{({\operatorname{mb}}),\ast}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, with number of bootstrap replications set to $B=1,000$. Recall that our results from the previous sections imply that the conditional bootstrap estimation errors $\hat{\mu}_{n}^{({\operatorname{cb}}(2)),*}-\hat{\mu}_{n}^{({\operatorname{cb}}(2))}$ and $\hat{\mu}_{n}^{({\operatorname{cb}}(3)),*}-\hat{\mu}_{n}^{({\operatorname{cb}}(3))}$ are consistent for the estimation error $\hat{\mu}_{n}^{({\operatorname{sb}})}-\mu_{r}$, that $\hat{\mu}_{n}^{({\operatorname{db}}),*}-\hat{\mu}_{n}^{({\operatorname{db}})}$ is consistent for $\hat{\mu}_{n}^{({\operatorname{db}})}-\mu_{r}$ and that $\hat{\mu}_{n}^{({\operatorname{sb}}),*}-\hat{\mu}_{n}^{({\operatorname{sb}})}$ is inconsistent for $\hat{\mu}_{n}^{({\operatorname{sb}})}-\mu_{r}$. All four statements are illustrated in Figure 3 by means of QQ-plots and histograms for the parameter choice $\gamma=-0.2,\beta=0.5$ and $m=80$; other choices lead to similar results.
Figure 3: Estimation error $\hat{\mu}_{n}^{({\operatorname{mb}})}-\mu_{r}$ vs. bootstrap error $\hat{\mu}_{n}^{({\operatorname{mb}}^{\prime}),*}-\hat{\mu}_{n}^{({\operatorname{mb}}^{\prime})}$ for mean estimators with fixed block size $r=90$. Top row: qq-plots. Bottom row: histograms and kernel density estimates. Underlying data from Model 6.1 with $\gamma=-0.2,\beta=0.5$ and $m=80$.
In the next step, we evaluate the performance of the four bootstrap approaches in terms of their ability to provide accurate estimates of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\mu}_{n}^{({\operatorname{mb}})})$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. For that purpose, for each fixed sample of size $n$, we estimate the respective variance by the empirical variance of the sample of bootstrap estimates; recall that each such sample is of size $B=1,000$, for every ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$. In Figure 4, we depict the average over the $N=5,000$ bootstrap estimates, along with the true parameters determined in a presimulation based on $10^{6}$ repetitions. We observe that the bootstrap estimates are close to the their target values for the disjoint method and for both circmax methods, while the naive sliding blocks bootstrap underestimates the true variance substantially, as mathematically demonstrated in Remark 4.3.
Figure 4: Bootstrap-based estimation of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\mu}_{n}^{({\operatorname{mb}})})$ with fixed sample size $n=1,000$. Left two columns: ${\operatorname{mb}}={\operatorname{sb}}$; the dashed line corresponds to $\sigma_{\operatorname{sb}}^{2}(r)$. Right two columns: ${\operatorname{mb}}={\operatorname{db}}$; the dashed line corresponds to $\sigma_{\operatorname{db}}^{2}(r)$.
Finally, we evaluate the performance of the bootstrap approaches in terms of their ability to provide accurate confidence intervals of pre-specified coverage; clearly, smaller intervals of similar coverage would be preferable. For that purpose, we restrict ourselves to the basic bootstrap confidence interval (Davison and Hinkley, , 1997) as illustrated in Corollary 5.6. The empirical coverage and the average widths of the respective intervals are depicted in Figure 5, where we omit the naive sliding method because of its inconsistency. We find that in most scenarios the desired coverage is almost reached by any method, in particular for lighter tail behavior. The disjoint blocks approach has the best coverage overall, albeit with only minimal advantages and sometimes on a par with or even slightly worse than the cb(2)-method. However, the price for this is universally wider confidence intervals.
Figure 5: Basic bootstrap confidence intervals for $\mu_{r}$ with fixed block size $r=90$. Top three rows: empirical coverage with intended coverage of $95\%$ (dashed line). Bottom three rows: relative average width with respect to the disjoint method, i.e., $\mathrm{width(CI}({\operatorname{db}}))/\mathrm{width(CI}({\operatorname{mb}}))$.
6.2 Fixed total sample size
In a second experiment, we consider the estimation of target parameters associated with the limiting law of $\bm{M}_{r}$ for $r\to\infty$, as well as bootstrap approximations for the respective estimation errors. For illustrative purposes, we restrict attention to the univariate, heavy tailed case and the estimation of the shape parameter $\alpha$ by pseudo-maximum likelihood estimation, as extensively discussed in Section 5.
Model 6.2 (ARMAX-Pareto-Model).
[…]
Performance of the estimators. In Figure 6 we depict the MSE of $\hat{\alpha}_{n}^{({\operatorname{mb}})}$ from Section 5 with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$ as a function of the effective sample size $m=n/r$. For $\beta=0.5$, we observe the typical bias-variance tradeoff resulting in a u-shaped MSE-curve: the larger the effective sample size, the smaller the variance and the larger the bias (for $\beta=0$, the MSE is dominated by the variance even for block sizes as small as 8). Apart from that, the findings are similar to the previous results for the case where $r$ was fixed: the sliding and circular block maxima estimator perform remarkably similar (in particular, the additional bias in the circular maxima is irrelevant), and they uniformly outperform the disjoint blocks estimator across all models under consideration.
Performance of the bootstrap. Next, we consider each of the bootstrap estimators $\hat{\alpha}_{n}^{({\operatorname{mb}}),\ast}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, with number of bootstrap replications set to $B=1,000$. Similar as in Section 6.1, we start by evaluating the performance of the four bootstrap approaches in terms of their ability to provide accurate estimates of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\alpha}_{n}^{({\operatorname{mb}})})$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. In Figure 7, we depict the average over the $N=5,000$ bootstrap estimates, along with the true parameters determined in a presimulation based on $10^{6}$ repetitions. The findings are akin to those in Section 6.1, the only additional remarkable observation being that the disjoint blocks method tends to overestimate the variance parameter in the iid case, in particular for large block sizes.
Figure 7: Bootstrap-based estimation of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\alpha}_{n}^{({\operatorname{mb}})})$ with fixed sample size $n=1,000$. Left two columns: ${\operatorname{mb}}={\operatorname{sb}}$; the dashed line corresponds to $\sigma_{\operatorname{sb}}^{2}(r)$. Right two columns: ${\operatorname{mb}}={\operatorname{db}}$; the dashed line corresponds to $\sigma_{\operatorname{db}}^{2}(r)$.
[…]
Figure 8: Basic bootstrap confidence intervals for $\alpha$ with fixed sample size $n=1,000$. Top three rows: empirical coverage with intended coverage of $95\%$ (dashed line). Bottom three rows: relative average width with respect to the disjoint method, i.e., $\mathrm{width(CI}({\operatorname{db}}))/\mathrm{width(CI}({\operatorname{mb}}))$.
7 Case study
[…]
Each moving window $w$ contains approximately $n=14,600$ daily observations, and the target parameter is estimated using the disjoint and sliding blocks estimator $\hat{\mu}_{n}^{({\operatorname{mb}})}=\hat{\mu}_{n}^{({\operatorname{mb}})}(w)$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. Basic bootstrap confidence intervals are constructed by either the classical approach based on resampling the disjoint block maxima (for estimator $\hat{\mu}_{n}^{({\operatorname{db}})}$), or by the circular block bootstrap approach with $k=2$ (for estimator $\hat{\mu}_{n}^{({\operatorname{sb}})}$). The resulting estimates and confidence bands are presented in Figure 9. The years on the x-axis correspond to the endpoints of the 40-year window. As expected from the theory (see in particular Figure 10), we find that the confidence intervals for the sliding blocks estimator are universally smaller than those for the disjoint blocks counterpart, with the average relative width being approximately 1.07.
Figure 9: Left plot: estimated expected annual maximum precipitation based on disjoint and sliding blocks estimators $\mu_{n}^{({\operatorname{db}})},\mu_{n}^{({\operatorname{sb}})}$ (solid lines), alongside with with $95\%$ bootstrap confidence intervals based on the db- and cb(2)-bootstrap, respectively (for a moving window of 40 years). Right plot: width of the respective confidence interval. All curves have been smoothed by a moving average filter of size 2.
8 Conclusion
Both the block-maxima method and the bootstrap are time-honored statistical methods that have seen wide use in applied statistics for extremes. Surprisingly, bootstrap consistency has never been proven, not even for the classical block-maxima method based on disjoint blocks. In this paper, respective consistency statements were established under high-level conditions on the data-generating process. A new approach, called the circular block-maxima method, has been proposed to allow for valid and computationally efficient bootstrap inference regarding the sliding block maxima method. The approach may be of independent interest for extreme-value analysis of non-stationary extremes, for instance in the presence of a temporal trend. Indeed, a reasonable model assumption inspired by many applications of the disjoint block maxima method would consist of the assumption that the circular block-maxima within a fixed $kr$-block of observations, say the $i$th one, follow an extreme-value distribution whose distribution/parameters are depending on $i$. Respective methods could be studied mathematically under a suitable triangular array structure, and we conjecture that advantages over the disjoint block maxima method will eventually show up in respective estimation variances. Other possible extensions would be the inclusion of inner-seasonal non-stationarities in the data-generating process or the generalization to empirical processes indexed by non-finite function classes.
9 Proofs
@@
Throughout, we omit the upper index $({\operatorname{cb}}).$ We start by showing that
$\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}\bm{h})=\Sigma_{\bm{h}}^{({\operatorname{sb}})}.$ (9.1)
@@
$\operatorname{Var}\big{(}\bar{\mathbb{G}}_{n,r}h\big{)}=k\operatorname{Var}(D_{n,1})+r_{n1}+r_{n2},$ (9.3)
where $r_{n1}=2k(1-\frac{1}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,2})$ and $r_{n2}=2k\sum_{d=2}^{m(k)-1}(1-\frac{d}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,1+d}).$ Thus, the proof of (9.1) is finished once we show that
$\lim_{n\to\infty}k\operatorname{Var}(D_{n,1})=\Sigma_{h}^{({\operatorname{sb}})},\quad\lim_{n\to\infty}r_{n1}=0,\quad\lim_{n\to\infty}r_{n2}=0.$ (9.4)
@@
where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|,(k)},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|,(k)})$ has cdf $G_{|\xi-\xi^{\prime}|}^{(k)}$ from (11.1) and where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|})$ has cdf $G_{|\xi-\xi^{\prime}|}$ from (2.3). Observe that, by Proposition 11.1, Condition 10.1 and Example 2.21 in van der Vaart, (1998), $\lim_{n\to\infty}f_{r}(\xi,\xi^{\prime})=g_{(k)}(\xi,\xi^{\prime})$. Hence, by Condition 10.1 and Dominated Convergence
@@
$=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+o(1).$ (9.5)
Now let $u(x):=g(x,0)$ and note that $g(\xi,\xi^{\prime})=u(|\xi-\xi^{\prime}|)$ and that $u(x)=0$ for $x>1$. By the definition of $G_{|\xi-\xi^{\prime}|}^{(k)}$ in (11.1) we have $G_{|\xi-\xi^{\prime}|}^{(k)}=G_{|\xi-\xi^{\prime}|}$ if $10\leq|\xi-\xi^{\prime}|\leq k-$ 1, which also implies that $g_{(k)}(\xi,\xi^{\prime})=g_{(k)}(|\xi-\xi^{\prime}|,0)=g(|\xi-\xi^{\prime}|,0)=u(|\xi-\xi^{\prime}|)$ for $\xi\leq k-1.$ Therefore,
@@
$=\int_{0}^{k}\int_{\xi}^{k\wedge(\xi+1)}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+\int_{0}^{1}\int_{k-1+\xi}^{k}g_{(k)}(\xi^{\prime}-\xi,0)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi,$ (9.6)
where we used that $u(x)=0$ for $x>1$ at the second equality. After substituting $x=\xi^{\prime}-\xi$ in the inner integral and noting that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2,$ the first integral on the right-hand side of (9) can be written as
@@
$=(k-1)\frac{\Sigma_{h}^{({\operatorname{sb}})}}{2}+\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y,$
where we used the substitution $y=k-\xi$ in the last step. Moreover, substituting $x=k-\xi^{\prime}+\xi$ in the inner integral, the second integral on the right-hand side of (9) can be written as
@@
where we used that $g_{(k)}(k-x,0)=u(x)$ for $x\in[0,1]$ by the definition of $G_{k-x}^{(k)}$. Overall, again using that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2$, the previous three displays imply that
$\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=k\frac{\Sigma_{h}^{({\operatorname{sb}})}}{2}.$
This implies (9.4) in view of the symmetry $\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{k}\int_{0}^{\xi}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ and (9).
Similar arguments as before, invoking asymptotic independence of $\bm{Z}_{r,s}$ and $\bm{Z}_{r,t}$ for $s\in I_{kr,1}$ and $t\in I_{kr,2}$ (see Proposition 11.1) and a similar identification of the covariances as integrals, imply that $r_{n1}=o(1)$. Finally, by Conditions 2.2(c) and 10.1 and Lemma 3.11 in Dehling and Philipp, (2002) we obtain
@@
since $2/\omega<\nu$, where we have used that there is a lag of at least $kr$ between the observations making up $D_{n,1}$ and $D_{n,1+d}$. Overall, we have shown all three assertions in (9.4), and hence the proof of (9.1) is finished.
To verify the asymptotic normality, by the Cramér-Wold Theorem, it is again sufficient consider the case $q=1$. Recalling that $\bar{\mathbb{G}}_{n,r}h$ is centered, the case $\Sigma_{h}^{({\operatorname{sb}})}=0$ follows from (9.1); thus assume $\Sigma_{h}^{({\operatorname{sb}})}>0.$ Following Bücher and Segers, 2018a , we use a suitable blocking technique. Choose $m^{\ast}=m_{n}^{\ast}\in\mathbb{N}$ with $3\leq m^{\ast}\leq m(k)$ such that $m^{\ast}\to\infty$ and $m^{\ast}=o((n/r)^{\nu/(2(1+\nu))})$ with $\nu$ from Condition 10.1. Next, define $p:=p_{n}:=m(k)/m^{\ast}$ and assume for simplicity that $p\in\mathbb{N},n/r\in\mathbb{N}$. Furthermore, let $J_{j}^{+}:=\{(j-1)m^{\ast}+1,\ldots,jm^{\ast}-1\},J_{j}^{-}:=\{jm^{\ast}\}$ and note that $|J_{j}^{+}|=(m^{\ast}-1).$ Then,
$\bar{\mathbb{G}}_{n,r}h=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}(S_{n,j}^{+}+S_{n,j}^{-}),\quad\text{ where }\quad S_{n,j}^{\pm}:=\sqrt{\frac{np}{r}}\frac{1}{m(k)}\sum_{i\in J_{j}^{\pm}}D_{n,i}$ (9.7)
with $D_{n,i}$ from (9.2). Noticing that the time series $(S_{n,j}^{-})_{j}$ is stationary, we have

### v2

@@ Appendix A Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
Appendix B Proofs
@@
The proof is a simplified version of our proof of Theorem 3.2 below, whence we omit further details. Under slightly different conditions, a proof can be found in Theorem B.1 in BucZan23 or Theorem 8.7 in BucSta24. ∎
Proof of Theorem 3.2.
Throughout, we omit the upper index $[{\operatorname{c}}].$ We start by showing that
$\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}\bm{h})=\Sigma_{\bm{h}}^{[{\operatorname{s}}]}.$ (B.1)
@@
$\operatorname{Var}\big{(}\bar{\mathbb{G}}_{n,r}h\big{)}=k\operatorname{Var}(D_{n,1})+r_{n1}+r_{n2},$ (B.3)
where $r_{n1}=2k(1-\frac{1}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,2})$ and $r_{n2}=2k\sum_{d=2}^{m(k)-1}(1-\frac{d}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,1+d}).$ Thus, the proof of (B.1) is finished once we show that
$\lim_{n\to\infty}k\operatorname{Var}(D_{n,1})=\Sigma_{h}^{[{\operatorname{s}}]},\quad\lim_{n\to\infty}r_{n1}=0,\quad\lim_{n\to\infty}r_{n2}=0.$ (B.4)
@@
where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|,(k)},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|,(k)})$ has CDF $G_{|\xi-\xi^{\prime}|}^{(k)}$ from (C.1) and where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|})$ has CDF $G_{|\xi-\xi^{\prime}|}$ from (2.3). Observe that, by Proposition C.1, Condition 9.1 and Example 2.21 in Van98, $\lim_{n\to\infty}f_{r}(\xi,\xi^{\prime})=g_{(k)}(\xi,\xi^{\prime})$. Hence, by Condition 9.1 and Dominated Convergence
@@
$=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+o(1).$ (B.5)
Now let $u(x):=g(x,0)$ and note that $g(\xi,\xi^{\prime})=u(|\xi-\xi^{\prime}|)$ and that $u(x)=0$ for $x>1$. By the definition of $G_{|\xi-\xi^{\prime}|}^{(k)}$ in (C.1) we have $G_{|\xi-\xi^{\prime}|}^{(k)}=G_{|\xi-\xi^{\prime}|}$ if $10\leq|\xi-\xi^{\prime}|\leq k-$ 1, which also implies that $g_{(k)}(\xi,\xi^{\prime})=g_{(k)}(|\xi-\xi^{\prime}|,0)=g(|\xi-\xi^{\prime}|,0)=u(|\xi-\xi^{\prime}|)$ for $\xi\leq k-1.$ Therefore,
@@
$=\int_{0}^{k}\int_{\xi}^{k\wedge(\xi+1)}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+\int_{0}^{1}\int_{k-1+\xi}^{k}g_{(k)}(\xi^{\prime}-\xi,0)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi,$ (B.6)
where we used that $u(x)=0$ for $x>1$ at the second equality. After substituting $x=\xi^{\prime}-\xi$ in the inner integral and noting that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{[{\operatorname{s}}]}/2,$ the first integral on the right-hand side of (B) can be written as
@@
$=(k-1)\frac{\Sigma_{h}^{[{\operatorname{s}}]}}{2}+\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y,$
where we used the substitution $y=k-\xi$ in the last step. Moreover, substituting $x=k-\xi^{\prime}+\xi$ in the inner integral, the second integral on the right-hand side of (B) can be written as
@@
where we used that $g_{(k)}(k-x,0)=u(x)$ for $x\in[0,1]$ by the definition of $G_{k-x}^{(k)}$. Overall, again using that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{[{\operatorname{s}}]}/2$, the previous three displays imply that
$\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=k\frac{\Sigma_{h}^{[{\operatorname{s}}]}}{2}.$
This implies (B.4) in view of the symmetry $\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{k}\int_{0}^{\xi}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ and (B).
Similar arguments as before, invoking asymptotic independence of $\bm{Z}_{r,s}$ and $\bm{Z}_{r,t}$ for $s\in I_{kr,1}$ and $t\in I_{kr,2}$ (see Proposition C.1) and a similar identification of the covariances as integrals, imply that $r_{n1}=o(1)$. Finally, by Conditions 2.2(c) and 9.1 and Lemma 3.11 in DehPhi02 we obtain
@@
since $2/\omega<\nu$, where we have used that there is a lag of at least $kr$ between the observations making up $D_{n,1}$ and $D_{n,1+d}$. Overall, we have shown all three assertions in (B.4), and hence the proof of (B.1) is finished.
To verify the asymptotic normality, by the Cramér-Wold Theorem, it is again sufficient consider the case $q=1$. Recalling that $\bar{\mathbb{G}}_{n,r}h$ is centered, the case $\Sigma_{h}^{[{\operatorname{s}}]}=0$ follows from (B.1); thus assume $\Sigma_{h}^{[{\operatorname{s}}]}>0.$ Following BucSeg18-sl, we use a suitable blocking technique. Choose $m^{\ast}=m_{n}^{\ast}\in\mathbb{N}$ with $3\leq m^{\ast}\leq m(k)$ such that $m^{\ast}\to\infty$ and $m^{\ast}=o((n/r)^{\nu/(2(1+\nu))})$ with $\nu$ from Condition 9.1. Next, define $p:=p_{n}:=m(k)/m^{\ast}$ and assume for simplicity that $p\in\mathbb{N},n/r\in\mathbb{N}$. Furthermore, let $J_{j}^{+}:=\{(j-1)m^{\ast}+1,\ldots,jm^{\ast}-1\},J_{j}^{-}:=\{jm^{\ast}\}$ and note that $|J_{j}^{+}|=(m^{\ast}-1).$ Then,
$\bar{\mathbb{G}}_{n,r}h=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}(S_{n,j}^{+}+S_{n,j}^{-}),\quad\text{ where }\quad S_{n,j}^{\pm}:=\sqrt{\frac{np}{r}}\frac{1}{m(k)}\sum_{i\in J_{j}^{\pm}}D_{n,i}$ (B.7)
with $D_{n,i}$ from (B.2). Noticing that the time series $(S_{n,j}^{-})_{j}$ is stationary, we have